Given the function F(s) = 8s + 7, find each of the following
F(1)=
F(-10)=
F(6.9)=

Answer by neatmath(302)


When dealing with functions, and values of functions, there is usually a very straightforward way to solve these types of questions.

In general, whenever you see the variable, in this case s, just replace the variable with the given number, in this case: 1, -10, and 6.9

F(s)=8s+7

F(1)=8*1+7=8+7=15

F(-10)=8*(-10)+7=(-80)+7=(-73)

F(6.9)=8*(6.9)+7=55.2+7=62.2

So we now know:

F(1)=15

F(-10)=(-73)

F(6.9)=62.25

I hope this helps!